# A two-minute burst of highly polarised radio emission originating from low Galactic latitude

### v1

@@ 5 Conclusions
While only a small number of ULPs have been discovered to-date, they are all confined to low galactic latitudes. The low galactic latitude of ASKAP J1755 combined with a lack of detections of similar bursts in the VAST survey points towards a population of variable sources confined to low Galactic latitudes and potentially the thin disk of the Milky Way. This discovery should motivate a comprehensive search for minute-timescale variability around the Galactic Centre and at low Galactic latitudes more generally, in order to characterise ULPs, other transient or variable Galactic sources, and potentially find other new classes of object.
@@ 6 Acknowledgments
Appendix A VAST Pipeline Search
The VAST pipeline (Pintaldi et al., 2022) takes images, noise maps and source catalogues from ASKAP observations, runs source association, carries out forced photometry where appropriate, and calculates source statistics. The resulting data can then be queried programmatically using vast-tools, a python package designed for interacting with VAST data (Stewart et al., 2023). For all searches we use the require the following criteria:
  • nearest neighbour distance larger than 1 arcminute;
  • number of relations equal to zero (i.e. the source is isolated).
A.1 Discovery
As part of VAST pipeline testing we ran a search for sources with a single detection on an early subset of the VAST Galactic survey including observations through to early 2023. We then used vast-tools (Stewart et al., 2023) to search for corresponding detections in circular polarisation. ASKAP J1755 was noted as a source of interest due to its particularly high flux density. We then carried out the more detailed analysis described in Section 2.2 and follow-up observations described in Section 2.3.
A.2 Subsequent Galactic search
We later performed a more comprehensive search using VAST-Galactic observations from 2022-11-14 to 2024-03-12 (1271 images in total), searching for sources similar to ASKAP J1755. There is reasonable coverage of each field, with the number of observations per field ranging from 26 to 33. We searched for all sources detected at most 3 times (i.e. a detection rate of $\lesssim 10\%$). After removing two clear artefacts and ASKAP J1755 we were left with seven new sources that passed our search criteria. Two sources have known multi-wavelength counterparts – one is the high-mass X-ray binary V* IL Lup and the other is a detection of Swift J151857.0-572147 previously reported by Anumarlapudi et al. (2024).
The four remaining sources have no clear multi-wavelength counterparts. After generating dynamic spectra using the procedure outlined in Appendix B we find that one is persistent across the 12 minute observation, while the others arise from short flares or bursts. The persistent source is unpolarised, but one flaring source exhibits circular polarisation, one exhibits linear polarisation and the third exhibits both. Two lie within $|b|\lesssim 1\degr$, while the others are within $|b|<4\degr$. In short, all share some properties with ASKAP J1755 but none are clear analogues based on the available data. Further investigation of these sources is ongoing and will be presented in future work.
A.3 Pilot search
We carried out a separate search of the VAST pilot survey, which consists of 1038 observations spanning 2019-04-25 to 2020-08-30. Full details of the survey can be found in (Murphy et al., 2021). We used similar criteria to above, with the only difference being relaxing the requirement of having 10 observations down to 5, reflecting the smaller survey size. The data used for this search has not been cropped to the central low-noise region, and hence we also removed sources that were outside the central $6.3\degr\times 6.3\degr$ square of the image footprint in order to remove image edge artefacts and make the search area consistent with that of the full survey. No sources passed qualitative inspection.
A.4 Extragalactic fields search
The VAST extragalactic survey began on 2023-06-13, with each field observed approximately every two months. We carried out a search of all observations through to 2024-05-09, consisting of 1850 images in total. Most fields have been observed 6–7 times and we therefore searched for sources with a single detection. After removing clear artefacts we were left with a single candidate that is associated with the star ASAS J055115+0355.3.
Appendix B Generation of dynamic spectra
We generated dynamic spectra and lightcurves using DStools [^7] https://github.com/joshoewahp/dstools/, a CASA based python package to construct and post-process dynamic spectra from calibrated visibilities, and describe the details of each processing stage below.
@@ Appendix B Generation of dynamic spectra
and constructed the total linear polarisation dynamic spectrum $L=\rm{Re}(Q)+i\rm{Re}(U)$ where $\rm{Re}(Q)$ and $\rm{Re}(U)$ are the real components of the Stokes $Q$ and $U$ dynamic spectra respectively, and $i$ is the imaginary unit. We performed Faraday rotation synthesis (Brentjens & de Bruyn, 2005b) using RMclean (Heald et al., 2009; Heald, 2017) to extract a rotation measure of $\rm{RM}=961\pm 45$ rad m-2 at the time of the burst peak, then corrected $L$ for Faraday rotation using
@@
Appendix C Inferred dispersion measure
@@
The first issue is not unexpected – a 10 s dispersion delay across the observed bandwidth corresponds to a dispersion measure of 2783 pc cm-3, which is in excess of the Milky Way contribution along this line of sight. To determine whether the recovered DM was sensible, we performed rigourous testing of this method by injecting Gaussian pulses into coarse-time-resolution white noise data with dispersion measures between 0-3000 $\rm{pc/cm^{3}}$. We considered two scenarios – a weak (S/N=5) pulse and a strong (S/N=20 and 50) pulse. Our simulation results showed a consistent measurement uncertainty of $\sim 200\,\rm{pc/cm^{3}}$, see Figure 7 for DM measurement results from weak pulses
@@
Figure 7: Recovered DM results from injection of weak pulses. We plot the offset between the measured value and the pulse DM on the y axis and the original injected pulse DM on the x axis.
@@
Appendix D Position measurement
@@
In order to better correct for the expected astrometric offsets and improve the overall astrometric uncertainty we performed an astrometric correction relative to the phase-referenced MeerKAT observations on a per-beam basis. We used Aegean (Hancock et al., 2012, 2018) to perform source-finding on the single-beam ASKAP image generated using the procedure in Section B and the deep MeerKAT image of the field generated using the procedure in Section 2.3.1. We filtered the list of sources from each image to exclude those with a peak flux density signal-to-noise ratio below 20, a compactness ratio (the ratio between the peak and integrated flux density) above 1.2 and those with a nearby source within $30\arcsec$ and $60\arcsec$ for MeerKAT and ASKAP respectively. This final filter removes source finding artefacts associated with bright sources, with the separation limits corresponding to approximately four beam-widths.

### v2

@@ 5 Conclusions
While only a small number of ULPs have been discovered to-date, they are all tightly confined to the Galactic plane unlike canonical radio pulsars. The low galactic latitude of ASKAP J1755 and the similarity of its polarisation properties with canonical pulsars should motivate a comprehensive search for minute-timescale variability around the Galactic Centre and at low Galactic latitudes more generally, in order to characterise ULPs, other transient or variable Galactic sources, and potentially find other new classes of object.
@@ 6 Acknowledgments
Appendix A Generation of dynamic spectra
We generated dynamic spectra and lightcurves using DStools (Pritchard, 2024), a CASA based python package to construct and post-process dynamic spectra from calibrated visibilities, and describe the details of each processing stage below.
@@ Appendix A Generation of dynamic spectra
and constructed the total linear polarisation dynamic spectrum $L=\rm{Re}(Q)+i\rm{Re}(U)$ where $\rm{Re}(Q)$ and $\rm{Re}(U)$ are the real components of the Stokes $Q$ and $U$ dynamic spectra respectively, and $i$ is the imaginary unit. We also de-biased the linearly polarised intensity using the method described by Müller et al. (2017). We performed Faraday rotation synthesis (Brentjens & de Bruyn, 2005b) using RMclean (Heald et al., 2009; Heald, 2017) to extract a rotation measure of $\rm{RM}=961\pm 45$ rad m-2 at the time of the burst peak, then corrected $L$ for Faraday rotation using
@@
Appendix B Inferred dispersion measure
@@
The first issue is not unexpected – a 10 s dispersion delay across the observed bandwidth corresponds to a dispersion measure of 2783 pc cm-3, which is in excess of the Milky Way contribution along this line of sight. To determine whether the recovered DM was sensible, we performed rigourous testing of this method by injecting Gaussian pulses into coarse-time-resolution white noise data with dispersion measures between 0-3000 $\rm{pc/cm^{3}}$. We considered two scenarios – a weak (S/N=5) pulse and a strong (S/N=50) pulse. Figure 8 shows the results of our simulations, where we recover the injected DM in both scenarios.
@@
Figure 8: Dispersion measure (DM) recovery for simulated pulses as described in Appendix B The injected pulse DM is shown on the x-axis, while the offset between the injected and measured DM values is shown on the y-axis. Left: DM recovery for a weak (S/N=5) pulse. Right: DM recovery for a strong (S/N=50) pulse.
Appendix C Position measurement
@@
In order to better correct for the expected astrometric offsets and improve the overall astrometric uncertainty we performed an astrometric correction relative to the phase-referenced MeerKAT observations on a per-beam basis. We used Aegean (Hancock et al., 2012, 2018) to perform source-finding on the single-beam ASKAP image generated using the procedure in Section A and the deep MeerKAT image of the field generated using the procedure in Section 2.3.1. We filtered the list of sources from each image to exclude those with a peak flux density signal-to-noise ratio below 20, a compactness ratio (the ratio between the peak and integrated flux density) above 1.2 and those with a nearby source within $30\arcsec$ and $60\arcsec$ for MeerKAT and ASKAP respectively. This final filter removes source finding artefacts associated with bright sources, with the separation limits corresponding to approximately four beam-widths.